2. A cash flow sequence starts in year 1 at $9000 and decreases by $250 each year through year 12. (a) Determine the value of the gradient G; (b) determine the amount of cash flow at the end of year 8; and (c) determine the value of n for the gradient.
Answer #1

a)

Gradient, G =Uniform annual change in cash flow=-$250

b)

If t is the number of year, and C1 is cash flow in year 1,

Cash flow at the end of year t =C1+(t-1)*G

Cash flow at the end of year 8=9000+(8-1)*(-250)=$7250

c)

Value of n=years in consideration=12 years